OpenStudy (anonymous):

Find the slope of the line whose equation is 5y = x - 3.

OpenStudy (anonymous):

slope= rise/run

OpenStudy (anonymous):

so rearrange for y

OpenStudy (anonymous):

hence y=x-3/5

OpenStudy (anonymous):

Thanks Again!

OpenStudy (anonymous):

so take to points on that line and let slope= y1-y2/x1-x2

OpenStudy (anonymous):

points being (x1,y1), and (x2,y2)